How many moles of NO2 are needed to increase the temperature of a 333-gram block of iron from 22⁰C to 1300⁰C? (specific heat of iron is 0.49 J/g⁰C)
3NO2 + H2O --> 2HNO3 + NO ∆H = - 138 kJ /mol
1
Expert's answer
2020-12-10T07:46:26-0500

First calculate the amount of heat needed to increase the temperature of a 333-gram block of iron from 22⁰C to 1300⁰C:

"Q = cm\Delta T = 0.49 * 333 * (1300-22) = 208531 (J) = 208.5 (kJ)"

If 3 moles of NO2 are needed to produce 138 kJ of heat, then calculate the amount of NO2 needed to produce 208.5 kJ of heat:

"\nu = 208.5*3/138 = 4.53 (mol)"

Answer: 4.53 moles of NO2 are needed to increase the temperature of a 333-gram block of iron from 22⁰C to 1300⁰C.